A simple pendulum uses a steel wire as the "string." The length of this wire is 1.5 m at room temperature. If the temperature is increased by 10°C, the length of the wire increases by 0.19 mm. What is the change in the period of the pendulum?
Answer #1

Here,

initial length , L = 1.5 m

change in length , x = 0.19 mm

x = 0.19 *10^-3 m

for the change in period is del T

for the change in period

del T /T = 0.5 * x/L

del T /T = 0.5 * 0.19 *10^-3/1.5

del T /T = 6.33 *10^-5

the period will charge by 6.33 *10^-5 times